# Polarization-Dependent Loss Mitigation with Orthogonal-Design-Based Precoding and Interference Cancellation

### v1

@@ Abstract
Recent work by Shehadeh and Kschischang provides a simple capacity-achieving scheme for channels with polarization-dependent loss (PDL) under common modeling assumptions via a careful choice of orthogonal-design-based precoding and interference cancellation. This letter extends that work with a simulation-based demonstration showing that this scheme remains highly effective at mitigating PDL in the highly practical setting of 4-PAM with Chase-decoded extended Hamming inner codes rather than the near-capacity inner codes considered in the original work. An alternative near-optimal variation of this scheme is provided requiring only one inner code rather than two and suffering no penalty in the absence of PDL, making it much more practical.
@@ I Setup
We further assume in this letter real-valued channels with real-valued inputs. However, we emphasize that everything presented generalizes to complex-valued channels and inputs by analogously considering the complex variant of the scheme included in [1]. We then have a two-parameter channel class
$\underbrace{\begin{bmatrix}Y_{1}\\ Y_{2}\end{bmatrix}}_{\mathbf{Y}}=\underbrace{\begin{bmatrix}\sqrt{1+\gamma}&0\\ 0&\sqrt{1-\gamma}\end{bmatrix}}_{\mathbf{D}_{\gamma}}\underbrace{\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix}}_{\mathbf{R}_{\theta}}\underbrace{\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}}_{\mathbf{X}}+\underbrace{\begin{bmatrix}Z_{1}\\ Z_{2}\end{bmatrix}}_{\mathbf{Z}}$
where $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$, representing an adversarial class of channels with two polarizations and up to
@@
of PDL where $\alpha\in[0,1)$ is a fixed parameter. Moreover, $\mathbf{X}$ and $\mathbf{Z}$ are independent with $\mathbf{Z}$ being standard white Gaussian and $\mathbf{X}$ satisfying $\mathbb{E}[\norm{\mathbf{X}}_{2}^{2}]=2\cdot\mathsf{SNR}$ where $\mathsf{SNR}$ denotes the signal-to-noise ratio (SNR). We are generally interested in the worst-case performance of any given scheme over all $\gamma\in[-\alpha,\alpha]$ and $\theta\in[0,2\pi)$.
When precoding, we consider a two-channel-use extension of the channel matrix given by $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\in\mathbb{R}^{4\times 4}$ which we can interpret either as representing two uses of the channel in time or the use of a complex-valued input with real-valued representations where the first and second halves of a vector represent the real and imaginary parts respectively. We then take the input to the channel to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{4\times 1}$ and $\mathbf{G}\in\mathbb{R}^{4\times 4}$ is an orthogonal precoding matrix. The effective channel is then described by the effective channel matrix $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\mathbf{G}$ and has input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=2\cdot\mathsf{SNR}$.
We use bit-interleaved coded modulation (BICM) [10] with extended Hamming codes and a Gray-labeled $4$-PAM constellation together with soft-decision Chase decoding [11]. Since the scheme of [1] requires two codes with different rates used in equal proportion, we choose them to have the same average rate as a third code which is used in schemes requiring one code. In particular, we consider three codes with lengths and dimensions $(n_{1},k_{1})$, $(n_{2},k_{2})$, and $(n,k)$ respectively satisfying
@@
We take $(n_{1},k_{1})=(100,90)$, $(n_{2},k_{2})=(400,388)$, and $(n,k)=(200,187)$. These codes are obtained by shortening longer extended Hamming parent codes. This yields a rate of $0.935$ so that the inner coding scheme is somewhat comparable to that in the 400ZR implementation agreement [12].
Our key contributions are the demonstration of the effectiveness of the scheme of [1] for PDL mitigation in this context and the introduction of a variation on this scheme with various practical advantages. We illustrate our points by comparing five schemes defined in Section II which include the scheme of [1] and the proposed variation. This is followed by simulations in Section III and concluding remarks in Section IV.
@@ II Five Schemes
We now describe five schemes which we identify as iZ, pZ, D, pD, and the scheme of [1]. These are based on precoding, linear minimum mean square error (LMMSE) equalization, and successive interference cancellation (SIC) [13, Chapter 8]. We use zero-forcing (ZF) instead of LMMSE for simplicity since it is comparable at the high SNRs of interest. Implementation details for all schemes can be found in our complete simulation code which we make available online [14].
@@ II-B One Code and ZF with Precoding (pZ)
This is followed by ZF and independent decoding of each codeword which sees a worst-case effective gain of $1-\alpha^{2}$. The resulting scheme is comparable to those of [15, 9].
@@ II-C One Code and D-BLAST-Style ZF-SIC without Precoding (D)
In Scheme D, we consider a D-BLAST-style [13, Chapter 8] approach which allows us to perform ZF-SIC on one codeword by diagonally staggering it in time and space. After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which is split into two halves $\mathbf{x}^{\mathsf{f}}$ and $\mathbf{x}^{\mathsf{s}}$ of $n/4$ symbols. We then transmit $\zeta$ such codewords in $(\zeta+1)n/4$ channel uses in time which we represent by the $42\times(\zeta+1)n/$ 4 matrix
@@
Figure 1: Simulation results for Schemes iZ, pZ, D, pD and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB.
@@ II-D One Code and D-Blast-Style ZF-SIC with Precoding (pD)
In Scheme pD, we consider combining the orthogonal-design-based precoding and interference cancellation approach of [1] with a D-BLAST-style technique as in Scheme D so that only a single code is used. In particular, we consider a two-channel-use extension and the precoding matrix (1) after which we apply Scheme D to the resulting effective channel. After mapping to a $4$-PAM constellation, we have a codeword of $n/2$ symbols which we split into two halves $\mathbf{u}^{\mathsf{f}}$ and $\mathbf{u}^{\mathsf{s}}$ of dimensions $82\times n/$ 8 containing $n/4$ symbols each. We then transmit $\zeta$ such codewords in $(\zeta+1)n/8$ uses of the two-channel-use extended channel so that the input is represented by the $84\times(\zeta+1)n/$ 8 matrix
@@ III Simulations
IV Concluding Remarks

### v2

# Polarization-Dependent Loss Mitigation via Orthogonal Design Precoding and Interference Cancellation

@@ Abstract
Recent work by Shehadeh and Kschischang provides a simple capacity-achieving scheme for channels with polarization-dependent loss (PDL) under common modeling assumptions via a careful choice of orthogonal-design-based precoding and interference cancellation. This letter extends that work with a simulation-based demonstration showing that this scheme remains highly effective at mitigating PDL in the highly practical setting of 16-QAM with Chase-decoded extended Hamming inner codes rather than the near-capacity inner codes considered in the original work. An alternative near-optimal variation of this scheme is also provided requiring only one inner code rather than two and suffering no penalty in the absence of PDL, making it much more practical.
@@ I Setup
By a standard reduction technique, the channel matrix under this model is given by
$\underbrace{\begin{pmatrix}\sqrt{1+\gamma}&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}}_{\mathbf{D}_{\gamma}}\underbrace{\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}}_{\mathbf{R}_{\theta}}\underbrace{\begin{pmatrix}e^{i\phi}&0\\ 0&e^{-i\phi}\end{pmatrix}}_{\mathbf{B}_{\phi}}$
which yields a three-parameter channel class
$\mathbf{Y}=\mathbf{D}_{\gamma}\mathbf{R}_{\theta}\mathbf{B}_{\phi}\mathbf{X}+\mathbf{Z}$
where $\gamma\in[-\alpha,\alpha]$, $\theta\in[0,2\pi)$, and $\phi\in[0,2\pi)$. This represents an adversarial class of channels corresponding to two polarizations with up to
@@
of PDL where $\alpha\in[0,1)$ is a fixed parameter. Moreover, $\mathbf{X}$ and $\mathbf{Z}$ are independent with $\mathbf{Z}$ being standard white Gaussian and $\mathbf{X}$ satisfying $\mathbb{E}[\norm{\mathbf{X}}_{2}^{2}]=2\cdot\mathsf{SNR}$ where $\mathsf{SNR}$ denotes the signal-to-noise ratio (SNR). We are generally interested in the worst-case performance of any given scheme over all $\gamma\in[-\alpha,\alpha]$, $\theta\in[0,2\pi)$, and $\phi\in[0,2\pi)$.
A common simplification whose justification is discussed in [2] is to take $\phi=0$ and thus remove $\mathbf{B}_{\phi}$ from the model. In this paper, we will cover both the simplified case of $\phi=0$ and the general case of adversarial $\phi\in[0,2\pi)$. In the case of $\phi=0$, it suffices to precode only across in-phase (I) and quadrature (Q) components and not across time, while in the general case of $\phi\in[0,2\pi)$, we must also precode across time. We will initially consider the case of $\phi=0$ for the purposes of exposition and because it is of interest in its own right, with the case of $\phi\in[0,2\pi)$ deferred to the end.
We will work with real-valued equivalent representations of complex-valued vectors and matrices where the first and second halves of the real-equivalent of a vector contain its real and imaginary parts respectively. This simplifies the description of I/Q precoding and the necessary widely-linear processing to equivalent linear processing on the concatenated real–imaginary representations. When $\phi=0$, the channel matrix is already real-valued and its real-equivalent representation is thus $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\in\mathbb{R}^{4\times 4}$. We take the input to the channel to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{4\times 1}$ and $\mathbf{G}\in\mathbb{R}^{4\times 4}$ is an orthogonal precoding matrix. The effective channel is then described by the $4\times 4$ effective channel matrix $\operatorname{diag}(\mathbf{D}_{\gamma}\mathbf{R}_{\theta},\mathbf{D}_{\gamma}\mathbf{R}_{\theta})\mathbf{G}$ and has input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=4\cdot\mathsf{SNR}$.
We use bit-interleaved coded modulation (BICM) [10] with extended Hamming codes and a $16$-QAM constellation. This corresponds to $\{-3,-1,1,3\}$-signaling in each degree of freedom under our real-equivalent representation. We further use soft-decision Chase decoding [11] in conjunction with all schemes considered. Since the scheme of [1] requires two codes with different rates used in equal proportion, we choose them to have the same average rate as a third code which is used in schemes requiring one code. In particular, we consider three codes with lengths and dimensions $(n_{1},k_{1})$, $(n_{2},k_{2})$, and $(n,k)$ respectively satisfying
@@
We take $(n_{1},k_{1})=(96,86)$, $(n_{2},k_{2})=(384,372)$, and $(n,k)=(192,179)$. These codes are obtained by shortening longer extended Hamming parent codes. This yields a rate of approximately $0.93$ so that the inner coded modulation scheme altogether is somewhat comparable to that in the 400ZR implementation agreement [12].
Our key contributions are the demonstration of the effectiveness of the scheme of [1] for PDL mitigation in this context and the introduction of a variation on this scheme with various practical advantages. We illustrate our points by comparing five schemes defined in Section II which include the scheme of [1] and the proposed variation followed by simulations in Section III. We then address more general scenarios in Section IV before concluding in Section V.
@@ II Five Schemes
We now describe five schemes which we identify as iZ, pZ, D, pD, and the scheme of [1]. These are based on precoding, linear minimum mean square error (LMMSE) equalization, and successive interference cancellation (SIC) [13, Chapter 8]. We use zero-forcing (ZF) instead of LMMSE for simplicity since it is comparable at the high SNRs of interest. We will assume $\phi=0$ and I/Q precoding as described in Section I for the purposes of describing our five schemes in this section. Implementation details for all schemes in both the $\phi=0$ case and the general $\phi\in[0,2\pi)$ case can be found in our complete simulation code which we make available online [14].
@@ II-B One Code and ZF with Precoding (pZ)
This is followed by ZF and independent decoding of each codeword which sees a worst-case effective gain of $1-\alpha^{2}$. The resulting scheme is essentially equivalent to those of [15, 9] as shown in [1]. In particular, I/Q and polarization–time coding are performance-equivalent when $\phi=0$.
@@ II-C One Code and D-BLAST-Style ZF-SIC without Precoding (D)
In Scheme D, we consider a D-BLAST-style [13, Chapter 8] approach which allows us to perform ZF-SIC on one codeword by diagonally staggering it in time and space. After Gray-mapping to $\{-3,-1,1,3\}$, we have a codeword of $n/2$ symbols which is split into two halves $\mathbf{x}^{\mathsf{f}}$ and $\mathbf{x}^{\mathsf{s}}$ of $n/4$ symbols. We then transmit $\zeta$ such codewords in $(\zeta+1)n/4$ channel uses in time which we represent by the $42\times(\zeta+1)n/$ 4 matrix
@@
Figure 1: Simulation results for Schemes iZ, pZ, D, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB.
@@ II-D One Code and D-Blast-Style ZF-SIC with Precoding (pD)
In Scheme pD, we consider combining the orthogonal-design-based precoding and interference cancellation approach of [1] with a D-BLAST-style technique as in Scheme D so that only a single code is used. In particular, we consider a two-channel-use extension and the precoding matrix (1) after which we apply Scheme D to the resulting effective channel. After Gray-mapping to $\{-3,-1,1,3\}$, we have a codeword of $n/2$ symbols which we split into two halves $\mathbf{u}^{\mathsf{f}}$ and $\mathbf{u}^{\mathsf{s}}$ of dimensions $82\times n/$ 8 containing $n/4$ symbols each. We then transmit $\zeta$ such codewords in $(\zeta+1)n/8$ uses of the two-channel-use extended channel so that the input is represented by the $84\times(\zeta+1)n/$ 8 matrix
@@ III Simulations
Figure 2: Simulation results for Schemes iZ, pZ, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.599$ corresponding respectively to worst-case PDLs of 00 dB and $6$ dB and in the case of a general strictly complex-valued channel.
@@
IV Further Results
When $\phi\in[0,2\pi)$, the channel matrix is complex-valued and has $4\times 4$ real-equivalent representation given by
[…]
which we extend to $\operatorname{diag}(\mathbf{H}_{\gamma,\theta,\phi},\mathbf{H}_{\gamma,\theta,\phi})\in\mathbb{R}^{8\times 8}$ representing two channel uses in time. We then take the input to be $\mathbf{X}=\mathbf{G}\mathbf{U}$ where $\mathbf{X},\mathbf{U}\in\mathbb{R}^{8\times 1}$ and $\mathbf{G}\in\mathbb{R}^{8\times 8}$ is our orthogonal precoding matrix. This yields an effective $8\times 8$ channel matrix given by $\operatorname{diag}(\mathbf{H}_{\gamma,\theta,\phi},\mathbf{H}_{\gamma,\theta,\phi})\mathbf{G}$ and having input $\mathbf{U}$ satisfying $\mathbb{E}[\norm{\mathbf{U}}_{2}^{2}]=8\cdot\mathsf{SNR}$. We then consider the $8\times 8$ precoding matrix from [1]
$\mathbf{G}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0&0&1&0&0&0\\ 0&0&1&0&0&0&1&0\\ 0&-1&0&0&0&-1&0&0\\ 0&0&0&-1&0&0&0&-1\\ 0&0&1&0&0&0&-1&0\\ -1&0&0&0&1&0&0&0\\ 0&0&0&1&0&0&0&-1\\ 0&-1&0&0&0&1&0&0\end{pmatrix}$ (2)
and straightforwardly apply the schemes of Section II with the effective channel split into two equal halves as before. The complete implementation details are available in our complete simulation code [14]. Simulation results are provided in Fig. 2 and are nearly identical to the case of $\phi=0$ as predicted by [1]. Scheme D is omitted since it was merely pedagogical.
Figure 3: Simulation results for Schemes iZ, pZ, pD, and the scheme of [1] for $\alpha=0$ and $\alpha=0.333$ corresponding respectively to worst-case PDLs of 00 dB and $3$ dB and in the case of a general strictly complex-valued channel.
Lastly, we consider the case of $3$ dB of worst-case PDL in Fig. 3. In this case, we change the parameters of the code pair used in the scheme of [1] to $(n_{1},k_{1})=(144,131)$ and $(n_{2},k_{2})=(288,275)$ which have the same average rate as before but are closer to each other to account for the smaller PDL. Note that the slightly worse performance of the scheme of [1] in this case relative to pD is not a contradiction of the information-theoretic optimality of the scheme of [1]. This is a finite blocklength effect due to the worse gap to capacity of the lower rate Hamming code in the code pair used. The scheme of [1] is always better if all inner codes used are at the same average gap to capacity. However, it is typically harder to achieve the same gap to capacity at lower rates that are still greater than $1/2$. This further highlights the significant benefits of the proposed scheme pD which obviates all of these issues and does not need to be adapted to the worst-case PDL value.
V Concluding Remarks